OpenStudy (anonymous):

sin(x+y)=y^2cosx dy/dx

OpenStudy (anonymous):

the derivative of function y, with respect to x. sin(x+y)=y^2cosx left side will have: cos(x+y) * (1+y' ) the right side will have: 2y * y' * cos x + y^2(-sin x) So we have altogether: cos(x+y) * (1+y' )=2y * y' * cos x + y^2(-sin x) now we just need to solve for y' . cos(x+y)+ y' *cos(x+y) = 2y * y' * cos x + y^2(-sin x) cos(x+y)+ y' *cos(x+y) = y' * 2ycos x + y^2(-sin x) - y' * 2y cosx + y' *cos(x+y) = -cos(x+y) + y^2(-sin x) y' = { y^2(-sin x) -cos(x+y) } / { cos(x +y) - 2y cos(x + y) }